### v2

@@ 5.3.2 Orthogonality with Model Updating Methods
LPIPS IN-C($\uparrow$) IN($\downarrow$) DDA 0.421 0.608 Decorruptor-DPM 0.575 0.573 Table 5: LPIPS scores with clean and corrupted images.
Avg. Acc (%) TENT 43.02 + Decorruptor-DPM 45.52 DeYO 48.61 + Decorruptor-DPM 49.50 Table 6: Orthogonality with model updates.
Decorruptor ResNet-50 Swin-T DPM (0.42s) 30.5 37.8 CM (1step) (0.05s) 26.8 34.5 4$\times$×CM (1step) (0.08s) 32.6 38.7 CM (4step) (0.10s) 27.5 35.6 4$\times$×CM (4step) (0.14s) 32.8 39.7 Table 7: Variants of Decorruptor-CM.
5.3.3 Analysis on Decorruptor-CM
Table 7 reports the experimental results for various choices of Decorruptor-CM. When editing the input image only once for prediction, it is fast but the performance decreases by about 2-3% compared to Decorruptor-DPM. However, obtaining four edited images for each input and using their average prediction in Eq. (10) leads to higher accuracy than DPM, regardless of the architecture. Notably, even for the same corrupted image, Decorruptor allows diverse edits towards different clean images. The increasing performance gap between 1 step and 4 steps as the model size grows suggests that the quality of images is inferred to be better with 4 steps.
5.4 Video Test-Time Adaptation
Decorruptor demonstrates significantly improved runtime compared to DDA. This indicates that Decorruptor can be practically applied not only in the image domain but also in the video domain. Therefore, we evaluated Decorruptor on a corrupted video dataset [32]. To align with the temporal property of videos, we applied Text2Video-Zero [26] to enable Decorruptor to generate temporally consistent frames. Text2Video-Zero utilizes cross-frame attention from the first frame across the entire sequence, enabling temporally coherent editing. As shown in Fig. 6, our Decorruptor-CM demonstrated superior performance in corruption editing compared to the DDA method. Moreover, for a 3-second input video, our Deccoruptor takes almost 10 seconds for corruption editing while DDA takes almost 20 minutes for corruption editing. Consequently, we demonstrate that our Decorruptor provides highly effective and efficient video corruption editing. More detailed quantitative results can be found in the Appendix.
6 Conclusion and Future Work
The existing diffusion-based image-level updating TTA approach is robust to data order and batch size variations but is impractical for real-world usage due to its slow processing speed. In response, we propose Decorruptor-DPM, leveraging a latent diffusion model for efficient memory and time utilization. Through fine-tuning via our novel corruption modeling scheme, Decorruptor-DPM possesses the capability to edit corrupted images. Additionally, we introduce Decorruptor-CM, employing consistency distillation to accelerate input updates further. Decorruptor surpasses the baseline diffusion-based approach in speed by 100 times while delivering superior performance. Furthermore, Decorruptor achieves a threefold increase in speed and superior performance compared to the data augmentation-based model updating baseline. However, Decorruptor does not completely address the issue observed in DDA, where it exhibits lower performance than the source-only model for certain corruptions. Addressing the challenge of consistently restoring any type of corruption to a clean image remains an ongoing task for image-level updating TTA approaches.
Appendix 0.A Additional Results
0.A.1 Detailed Results of Image Corruption Editing
Tables 8 and 9 present detailed performance results of Decorruptor on ImageNet-C [15] and ImageNet-$\bar{\mathrm{C}}$ [39], respectively, using ResNet-50 [14] as the architecture. As shown in Table 1, Decorruptor demonstrates significant performance improvement for Noise and Weather corruptions. However, for some of Digital corruptions, it exhibits slightly lower performance than source-only for some specific corruptions. DDA [13] also shares the limitation of not being able to properly edit some corruptions. Addressing this issue is crucial for the effectiveness of the input updating TTA [59] method. As illustrated in Table 2, Our Decorruptor shows consistent improvement for all corruptions in ImageNet-$\bar{\mathrm{C}}$. Commonly, ensembling more edited images always presents performance improvement for all corruptions in ImageNet-C and ImageNet-$\bar{\mathrm{C}}$. Further results of ensembling is provided in Section 0.A.5.
@@ 0.A.1 Detailed Results of Image Corruption Editing
0.A.2 Quantitative Video Corruption Editing Results
For our experiments, we referred to the performance chart of ViTTA (see Table. 2 in Lin et al. [32]). In this section, we describe the results obtained by applying Decorruptor-CM for video corruption editing. The UCF-101C [32] dataset includes 3,783 corrupted videos for each type of corruption, totaling 10 different corruptions. The entire process of video decorruption was conducted using eight A40 GPUs and took about three days. The network used in the experiments was TANet [34], and when combining the model update method with our approach, we did not use an ensemble but instead assessed the performance solely using the generated dataset. The results are described in Table. 10. As a result, when ensembling with the source, there was an average performance improvement of approximately 13% compared to source-only inference. These results suggest that our Decorruptor-CM can be effectively applied to video domains. Moreover, by applying our method with the TTA methodology, we confirmed an average performance improvement of about 3% compared to ViTTA, especially showing robust decorruption results against noise.
@@ 0.A.2 Quantitative Video Corruption Editing Results
Table 10: Qualitative results for the UCF101-C dataset. Here, ‘Ours-Only’ refers to results obtained from inference using only input updates. We further provide the results of combining our methodology with the baseline TTA method.
0.A.3 Corruption Granularity
Our proposed Decorruptor-DPM and CM methodologies also exhibit superior decorruption capabilities across all levels of severity when compared with DDA [13]. Notably, as depicted in Fig. 7 (b), CM shows comparable results with DPM only with 4 NFEs while effectively preserving the object-centric regions of a given image. Note, background colors sometimes change due to the stochastic nature of the diffusion model.
@@ 0.A.4 Further Use-Cases
0.A.5 Further Results of Ensemble
@@
As shown in Table 7 of Section 5.3 in the main manuscript, the addition of ensemble in Decorruptor-CM led to performance improvement. However, without careful consideration, increasing the number of edited images required for ensemble results in drawbacks in terms of runtime and memory consumption. Therefore, the number of edited images also becomes a crucial hyperparameter. We illustrated the performance variations with the change in the number of images used for ensemble in ImageNet-C and ImageNet-$\bar{\mathrm{C}}$ in Figures 9 and 10, respectively. The results indicate a consistent performance increase regardless of the architecture as the number of edited images increases, and the performance tends to converge around 4 ensembles.
@@
Figure 9: The accuracy (%) according to the number of edited images for ensembling in ImageNet-C
@@
Figure 10: The accuracy (%) according to the number of edited images for ensembling in ImageNet-$\bar{\mathrm{C}}$
0.A.6 Diverse Corruption Scenarios, Image Sizes, and Domains
@@
In this section, we presume a realistic situation when an image with various mixed corruptions is incoming at test time. We validate the editing performance for this scenario. For this experiment, we use Decorruptor-CM with 4 NFEs. As a result, as described in Fig. 11, we confirm that corruption editing is feasible with mixed corruption scenarios not only for (a) in-domain images but also for (b) out-of-domain images (e.g., panorama images). It is worth noting that we use mixed corruption severity as 5 for each corruption. Note that in each figure, the left side presents the corrupted images, while the right side displays the edited counterparts.
@@
Figure 11: Visualization of experimental results on (a) in-domain, and (b) out-of-domain corruption editing performance. We confirmed that our proposed method robustly performs corruption editing even in scenarios where mixed corruption is presented at test time.
Appendix 0.B Ablation Studies
0.B.1 Image Guidance Scaling on Consistency Model
@@
Figure 12: Ablation studies on (a) using the image guidance scale as conditioning, (b) fixed image guidance scale as 1.3, and (c) not using image guidance scale conditioning while distillation.
We describe the advantages of using the image guidance scale learnable in CM. LCM [37] shows fast convergence and a significant performance boost in distillation by using learnable guidance scales on SD. To extend LCM of SD to IP2P pipeline, we design a new learnable image guidance scale conditioning and enable multi-modal (i.e., image and text) conditioning.
@@
For each ablation experiment in Fig. 12, we trained Decorruptor-CM for 12K iterations, consuming 24 GPU hours on an A40 GPU, respectively. As a result, as seen in (a), applying the proposed method by combining it with a text-guidance scale demonstrated the highest performance on corruption editing compared to others. For experiments, we use corruptions such as checkerboard, Brownian noise, and Gaussian noise. In the following, when the image guidance scale scheduling is not used for distillation, as shown in (b), we observe that weird images are generated. In the following in (c), when the guidance scale is fixed for distillation, editing is hardly performed, remaining close to the given image semantics. Here, the image guidance scale used is 1.3. Therefore, we confirm the importance of learnable image guidance scale conditioning on the distillation stage for effective corruption editing.
0.B.2 Using Other Fast Diffusion Schedulers
@@
We conducted experiments based on the type of scheduler using the DPM-Solver++ sampler [36], traditionally utilized for fast sampling. The results, as shown in Fig. 13 (a), indicate that the sample quality of edited images dramatically decreases with smaller NFEs, and catastrophic failure occurs at 1 NFE. On the other hand, intriguingly, as shown in Fig, 13 (b), our Decorruptor-CM showed comparable corruption editing performance even at 1 NFE to that at 4 NFEs and confirmed better editing quality as the NFE increases. Therefore, this suggests that our proposed Decorruptor-CM enables fast, high-performance corruption editing. Each experiment was conducted with a fixed random seed.
Appendix 0.C Failure Cases
@@
Although our method consistently outperforms other baselines, noticeable improvements were not observed in editing like blur, and pixelation corruptions. Furthermore, as illustrated in Fig. 14, when presented with more realistic degradations such as fog or raindrops, we have observed that our model exhibits limitations in corruption editing. Although containing paired datasets in the pre-training stage can resolve such realistic degradations, methods of how to edit such images at test-time without containing such realistic paired images (i.e., clean and corrupted) during training remain a challenging problem for TTA researchers.

### v3

@@ 0.B.1 Detailed Results of Image Corruption Editing
0.B.2 Comparisons with EM-Based TTA Methods
In Table 12, we present additional comparisons of our computational costs in terms of accuracy and memory compared to existing EM-based TTA methods. As a result, adding Decorruptor-4$\times$×CM resulted in significant improvements across all approaches and datasets. In terms of efficiency, DDA incurs an additional runtime of 19.5s, whereas 4$\times$×CM adds an additional runtime of 0.14s, which is only about three times the runtime of EM-based methods (0.05s). Decorruptor shows significant improvements compared to the state-of-the-art EM-based method DeYO (48.6% $\rightarrow$ 51.6%) and can be applied to other downstream tasks (e.g., VideoTTA), demonstrating its strong superiority.
Table 12: Comparisons with TTA methods.
Source + DPM + 4$\times$×CM EATA + DPM + 4$\times$×CM SAR + DPM + 4$\times$×CM
Runtime (s/sample) 0.004 + 0.42 + 0.14 0.047 + 0.42 + 0.14 0.054 + 0.42 + 0.14
Memory (MB) 2,340 + 4,602 + 4,958 2,704 + 4,602 + 4,958 2,702 + 4,602 + 4,958
IN-C acc (%) 18.0 31.2 33.8 47.8 47.5 51.6 44.0 47.4 49.6
IN-$\bar{\mathrm{C}}$ acc (%) 25.0 41.6 47.7 54.0 57.6 59.7 49.9 55.9 58.8
0.B.3 Quantitative Video Corruption Editing Results
In this section, we elaborate on the results obtained by applying Decorruptor-CM for video corruption editing. As shown in Fig. 6, Decorruptor-CM outperformed DDA in corruption editing. For a 3-second input video, Decorruptor takes about 10 seconds, while DDA takes nearly 20 minutes. This demonstrates that Decorruptor is both highly effective and efficient for video corruption editing. For our experiments, we referred to the performance chart of ViTTA (see Table 2 in Lin et al. [34]). The UCF-101C [34] dataset includes 3,783 corrupted videos for each type of corruption, covering a total of 12 different corruptions. The entire process of video decorruption was conducted using eight A40 GPUs and took about three days. The network used in the experiments was TANet [36]. Instead of using an ensemble, we assessed the performance solely using the generated dataset when combining the model update method with our approach. The results are described in Table 13. Ensembling with the source resulted in an average performance improvement of approximately 13% compared to source-only inference. These findings suggest that our Decorruptor-CM can be effectively applied to video domains. Additionally, by applying our method with the TTA methodology, we observed an average performance improvement of about 3% compared to ViTTA, particularly showing robust decorruption results against noise.
Table 13: Quantitative results for the UCF101-C dataset. Here, ‘Ours-Only’ refers to results obtained from inference using only input updates. We further provide the results of combining our methodology with the baseline TTA method.
@@ 0.B.3 Quantitative Video Corruption Editing Results
0.B.4 Corruption Granularity
Our proposed Decorruptor-DPM and CM methodologies also exhibit superior decorruption capabilities across all levels of severity when compared with DDA [14]. Notably, as depicted in Fig. 7 (b), CM shows comparable results with DPM only with 4 NFEs while effectively preserving the object-centric regions of a given image. Note that background colors sometimes change due to the stochastic nature of the diffusion model.
@@ 0.B.5 Further Use-Cases
0.B.6 Additional Results of the Ensemble
@@
As shown in Table 3 of Section 5.2 in the main text, the addition of an ensemble in Decorruptor-CM led to performance improvement. However, without careful consideration, increasing the number of edited images required for an ensemble results in drawbacks in terms of runtime and memory consumption. Therefore, the number of edited images also becomes a crucial hyperparameter. We illustrated the performance variations with the change in the number of images used for the ensemble in ImageNet-C and ImageNet-$\bar{\mathrm{C}}$ in Figures 9 and 10, respectively. The results indicate a consistent performance increase regardless of the architecture as the number of edited images increases, and the performance tends to converge around 4 ensembles.
@@
Figure 9: The accuracy (%) according to the number of edited images for ensembling in ImageNet-C.
@@
Figure 10: The accuracy (%) according to the number of edited images for ensembling in ImageNet-$\bar{\mathrm{C}}$.
0.B.7 Diverse Corruption Scenarios, Image Sizes, and Domains
@@
In this section, we consider a realistic scenario where an image with various mixed corruptions is encountered at test time. We evaluate the editing performance for this situation using Decorruptor-CM with 4 NFEs. As shown in Fig. 11, we confirm that corruption editing is feasible in mixed corruption scenarios for both (a) in-domain images and (b) out-of-domain images (e.g., panorama images). We use a mixed corruption severity level of 5 for each type of corruption. In each figure, the left side presents the corrupted images, while the right side displays the edited counterparts.
@@
Figure 11: Visualization of experimental results on (a) in-domain and (b) out-of-domain corruption editing performance. We confirmed that our proposed method robustly performs corruption editing even in scenarios with mixed corruption at test time.
Appendix 0.C Ablation Studies
0.C.1 Image Guidance Scaling on Consistency Model
@@
Figure 12: Ablation studies on (a) using the image guidance scale as conditioning, (b) fixed image guidance scale as 1.3, and (c) not using image guidance scale conditioning during distillation.
@@
For the ablation study, we trained Decorruptor-CM under three conditions: (a) using our multi-modal guidance, (b) using a fixed image guidance, and (c) without using image guidance. Each experiment involved training the model for 12K iterations, consuming 24 GPU hours on an A40 GPU. Fig. 12 highlights the importance of our multi-modal guidance scale. We demonstrate corruption editing performance for checkerboard, Brownian noise, and Gaussian noise. As seen in (a), applying the proposed method by combining it with a text-guidance scale demonstrated the highest performance in corruption editing. In (b), we observe that abnormal images are generated when image guidance scale scheduling is not used for distillation. In (c), editing is minimal when the guidance scale is fixed during distillation, with the images remaining close to the original semantics. Thus, we confirm the importance of a learnable image guidance scale during the distillation stage for effective corruption editing. It is worth noting that guidance scheduling is considered not in CM inference, but only in DPM inference.
0.C.2 Using Other Fast Diffusion Schedulers
@@
We conducted experiments based on the type of scheduler using the DPM-Solver++ sampler [38], traditionally utilized for fast sampling. The results, as shown in Fig.13 (a), indicate that the sample quality of edited images dramatically decreases with smaller NFEs, with catastrophic failure occurring at 1 NFE. Conversely, as shown in Fig.13 (b), our Decorruptor-CM demonstrated comparable corruption editing performance at 1 NFE to that at 4 NFEs and showed better editing quality as the NFE increased. This suggests that our proposed Decorruptor-CM enables fast, high-performance corruption editing. Each experiment was conducted with a fixed random seed.
Appendix 0.D Failure Cases
@@
Although our method consistently outperforms other baselines, noticeable improvements were not observed in editing blur and pixelation corruptions. Furthermore, as illustrated in Fig. 14, our model exhibits limitations in corruption editing when faced with more realistic degradations, such as fog or raindrops, which could not be accurately modeled in our corruption modeling scheme. While including paired datasets in the pre-training stage can address such realistic degradations, finding methods to edit such images at test time without having realistic paired images (i.e., clean and corrupted) during training remains a challenging problem for TTA researchers.